# Exploring Precision and Recall to assess the quality and diversity of LLMs

### v2

@@ Abstract
The implementation of CAME is publicly available This paper introduces a novel evaluation framework for Large Language Models (LLMs) such as Llama-2 and Mistral, focusing on the adaptation of Precision and Recall metrics from image generation to text generation. This approach allows for a nuanced assessment of the quality and diversity of generated text without the need for aligned corpora. By conducting a comprehensive evaluation of state-of-the-art language models, the study reveals significant insights into their performance on open-ended generation tasks, which are not adequately captured by traditional benchmarks. The findings highlight a trade-off between the quality and diversity of generated samples, particularly when models are fine-tuned with human feedback. This work extends the toolkit for distribution-based NLP evaluation, offering insights into the practical capabilities and challenges that current LLMs face in generating diverse and high-quality text. We release our code and data [^1] https://github.com/AlexVerine/pr-4-llm.
@@
[…]
@@ 1 Introduction
In recent years and months, there has been a rapid democratization of Large Language Models (LLMs), exemplified by platforms such as ChatGPT and HuggingChat. These models are now widely accessible for a diverse range of tasks, from composing emails and application letters to performing multi-document summarization, generating medical prescriptions, and even crafting poetry and novels. The expanding spectrum of applications underscores the ubiquity of LLMs in our daily lives, necessitating the development of novel evaluation frameworks to accommodate their growing relevance.
Figure 1: Precision and Recall of various models on generating the WebText dataset, with the 2 standard deviation error ellipsis. Chat and pre-trained models different behaviors are clearly captured by our metrics.
@@
In contrast, the recent line of work on distribution-based metrics greatly departs from the sampled based evaluation. By considering LLMs and datasets as empirical distributions, the new metrics attempt to quantify how they differ and how they overlap. This shift drastically changes the scope of the evaluation. Beyond performance measures based on human references, the goal is to estimate the disparity between some data distribution ($P$) exhibited by human-written texts and the distribution learned by a LLM ($Q$), eliminating the need for aligned corpora. As exemplified by the development of MAUVE Pillutla et al. (2021), this kind of approach opens new perspectives to really compare LLMs in terms of their generative abilities rather than for some peculiar tasks.
However, many factors take part in assessing open-ended text generation and focusing on a single measure may restrict the significance of the evaluation. This is well known in the field of Information Retrieval, for example, where Precision and Recall are at the core of evaluation. While they have been previously tailored for image generation by Sajjadi et al. (2018) and Kynkäänniemi et al. (2019), we propose to adapt them to LLMs. With an extensive set of experiments involving prominent LLMs such as Llama-2 and Mistral, we show that these two new metrics significantly improve the evaluation with better understanding and characterization of the flaws in text generation. Precision and Recall allow us to clearly distinguish between samples quality or adequacy (for Precision) and a lack of diversity in the model outputs (for Recall). Empirical results show that these two measures are necessary for in-depth comparison of LLMs. For instance, we are able to quantify that fine-tuning models with human feedback significantly improves sample quality, albeit at the expense of sample diversity, as evidenced in Figure 1 by the trade-off between Precision and Recall.
@@
  • Our investigation sheds light on the impact of fine-tuning LLMs with human feedback. We present empirical evidence that, while such fine-tuning can improve the quality of generated samples, it also reduce their diversity, highlighting a crucial trade-off between Precision and Recall.
@@
Figure 2: Illustration of what can bring Precision and Recall compared to MAUVE. We consider a reference dataset $P$ composed of articles from 2 labels, World and Sport. $Q_{2}$ is made of articles from the same distribution. We compare it with two other datasets: $Q_{1}$ composed only of World articles and $Q_{3}$ composed of even numbers of World, Sport, Business and Sci/Tech articles. Relatively to $Q_{2}$, the MAUVE scores of $Q_{1}$ and $Q_{3}$ are almost identical, while Precision and Recall help differentiating how the distributions actually differ from the reference $P$.
@@ 2 Related works
Historically, assessing text generation tasks has been dependent on human resources, such as annotated texts or aligned corpora. This is particularly true in areas like machine translation and automatic summarization, where N-gram based metrics have been developed, notably BLEU (Papineni et al., 2001) and ROUGE (Lin, 2004). These metrics perform a rough comparison between the generated text and human-produced and aligned reference texts. More recent developments have introduced model-based metrics, such as BERTScore [^2] BERTScore uses precision and recall scores that differ significantly from ours, as they evaluate the similarity between a single generated text and its aligned reference (Zhang* et al., 2020), which aim to capture semantic similarities more effectively. However, these evaluations still largely depend on "aligned references," meaning human-provided examples that constrain the expected range of generated outputs. The challenge becomes significantly greater with open-ended generation tasks, where the range of acceptable responses is vast and cannot be encapsulated by a single reference.
@@ 2.1 Standard evaluation of generative models
Perplexity is the historical measure for language modeling Bahl et al. (1977). Easy to use and cheap to compute, practical motivations clearly explain its persistency for evaluation, even if it has notable limitations. As demonstrated in the work of Pillutla et al. (2021), perplexity works at the token level, only considering the surface forms, hence missing semantic information. Moreover, the metric is calculated on a per-sample basis, thereby failing to capture the diversity of generated texts.
Nowadays, the prevalent method for evaluating recent LLMs involves benchmarks with many tasks with easily verifiable correct answers (Touvron et al., 2023; Google, 2023; Beeching et al., 2023). Typically, this entails the generation of responses to short questions prompted in the instructions. The responses are then compared to a reference using metrics such as Exact-Match. While effective in assessing the comprehension and reasoning capabilities of LLMs, this kind of approach does not assess the generation skills of the models.
To overcome the limitations of automatic metrics, another trend considers human evaluation, especially for complex tasks. Beyond its cost, the guidelines must be carefully designed, since various criteria could be necessary to rate texts and compare model outputs. At the end human evaluation is prone to high variance and is difficult to reproduce. Another issue, is that diversity is not evaluated is standard human evaluation protocols, since texts are rated individually or by pairs.
@@ 2.2 Evaluation of open-ended generation
While the prevailing evaluations primarily focus on the quality of generated texts, some studies aim to characterize the diversity of these outputs. One approach involves computing the proportion of distinct N-grams generated, as demonstrated in Li et al. (2016). Another notable proposition by Zhu et al. (2018) introduces Self-BLEU which computes the BLEU score of each generated text against all other generated texts as a reference. However, N-gram matching cannot capture semantic features. For instance, a model generating random words achieves a perfect Self-BLEU score.
In the current landscape, most of the existing methods either prioritize quality or rely on simplistic characterization of diversity. Nevertheless, the recent introduction of MAUVE Pillutla et al. (2021) makes a significant shift toward diversity. By drawing inspiration from the domain of image generation, this new metric compares generated texts to a reference distribution without requiring aligned texts. MAUVE consists of a divergence-based metric that captures certain properties of text and exhibits correlation with human judgment. Subsequent research by Pimentel et al. (2023) confirms the validity of this approach by leveraging common divergence measures. MAUVE and subsequent works propose a simple metric that allows summarizing the evaluation of both quality and diversity in a single measure. Our work builds upon these efforts; we aim to demonstrate that using two distinct measures enables to distinguish between lack of quality and lack of diversity. This distinction becomes necessary to enable a deeper and more precise understanding of generative models, as illustrated in a simple example in Figure 2.
@@
Figure 3: Example of distribution of images. $P$ is the reference distribution of images of the CelebA dataset Liu et al. (2015), $Q_{1}$ and $Q_{2}$ are two different distributions of images. $Q_{1}$ has high quality, but low diversity, while $Q_{2}$ has high diversity and a low quality. Numbers and images are from Kynkäänniemi et al. (2019).
@@ 3 Background on Precision and Recall
Nowadays generative models for images are able to output ultra-realistic samples, making classical metrics such as Heusel et al. (2017) too coarse to assess the full distributional properties of the models. This jump of performance has motivated the definition of precision and diversity for generative models. This is exemplified on Figure 3. $Q_{1}$ represents a model that produces realistic images but lacks diversity in the generated people’s appearance. Conversely, $Q_{2}$ depicts a model generating a broad range of people with lower quality. Capturing this significant tradeoff is essential for comprehending generative model behavior. Inspired by the precision and recall metrics of binary classifications, the pioneering work of Sajjadi et al. (2018) proposed to redefine these two terms to assess the quality and diversity of generative models. These metrics are based on a discretized estimation of the reference and model’s distribution. Kynkäänniemi et al. (2019) pointed out some limitations of those metrics and proposed a new way to compute the Precision and Recall, leveraging more robust support estimation. This method is now widely adopted in the literature and is typically used along with the Fréchet Inception Distance Heusel et al. (2017). Still, these metrics have some limitations, like the sensibility to outliers (Naeem et al., 2020). This motivates further ongoing work on this topic, including theoretical contributions (Simon et al., 2019; Naeem et al., 2020; Alaa et al., 2022; Cheema and Urner, 2023; Verine et al., 2023) or practical methods (Kim et al., 2023).
@@
Given its importance in the field of image generation, we adopt the definition of Precision and Recall introduced by Kynkäänniemi et al. (2019). Alternative definitions and approaches are discussed in Appendix A.
@@
Since Precision and Recall involve an estimation of the support, standard pipelines involve a projection of the samples into an latent space, using pre-trained models such as Inception-v3 Szegedy et al. (2015) or VGG Simonyan and Zisserman (2015). Support of the distributions is then estimated using a $k$-nearest neighbors algorithm. We present this algorithm in more detail in Section 4.
@@
Although the practice of reporting Precision and Recall is well-established in image generation Dhariwal and Nichol (2021); Sauer et al. (2022); Song et al. (2023), it is a very recent practice in the field of text generation. Pillutla et al. (2021) led the way in this field by introducing MAUVE, a pioneering metric that builds on the theoretical concepts of quality and diversity, as initially proposed by Djolonga et al. (2020), and detailed in Appendix A.
The authors showcased the intriguing characteristics of their metrics across various GPT2 models and decoding algorithms. This metric is highly effective in evaluating how well the generated distribution aligns with the reference. The methodology behind it was subsequently validated and simplified by Pimentel et al. (2023) and further evaluated with alternative frameworks by Pillutla et al. (2023).
Despite the progress made, these approaches present a challenge: interpreting them as a measure of quality and diversity is not straightforward as they condense these two concepts into a single divergence measure. Characterizing Precision and Recall independently has only been mention by Pillutla et al. (2023) who conducted some toys experiments on small models or decoding algorithms.
Therefore, we believe that Precision and Recall for text generation is in its early stages. We argue that a deeper understanding of these metrics and their practical implications is needed.
@@ 4 Precision and Recall to assess text generation
In this paper, we introduce a framework for calculating Precision and Recall between distributions on texts. These distributions characterize a dataset, or a model and its generative capability. The extended version of Precision and Recall builds upon the notion of support of distributions, which is in our case challenging to define and estimate.
@@ 4.1 Pipeline to compute Precision and Recall
    Select $N$ sequences randomly in the dataset to build the set of references: $\mathcal{X}^{\mathrm{ref}}=\left\{{\bm{x}}_{1}^{\mathrm{ref}},\dots,{\bm{x}}_{N}^{\mathrm{ref}}\right\}.$ Sample $N$ sequences from the distribution $Q$ to assess and build the set of outputs: $\mathcal{X}^{\mathrm{out}}=\left\{{\bm{x}}_{1}^{\mathrm{out}},\dots,{\bm{x}}_{N}^{\mathrm{out}}\right\}.$
@@
    Precision and Recall computation:
@@
To choose the other parameters – the number of samples $N$ and the neighbor parameter $k$ – we investigate how Precision and Recall are affected by their variations. We utilize the WebText generation task with Llama-2 7B Chat (comprehensively described in Section 5.3) as a test case, prompting the model to generate between 100 to 10,000 samples. The findings are depicted in Figure 6. Additional details can be found in Section B.1.
Our observations are: (1) Precision and Recall values stabilize with an increasing $N$, plateauing at approximately $N=3000$. (2) As $k$ increases, Precision and Recall values gradually approach 1, aligning with the expectation that a larger $k$ leads to a wider estimated support. In the rest of the paper, we set $N$ to $4000$ samples and $k$ to $4$.

### v3

@@ Abstract
We introduce a novel evaluation framework for Large Language Models (LLMs) such as Llama-2 and Mistral, focusing on importing Precision and Recall metrics from image generation to text generation. This approach allows for a nuanced assessment of the quality and diversity of generated text without the need for aligned corpora. By conducting a comprehensive evaluation of state-of-the-art language models, the study reveals new insights into their performance on open-ended generation tasks, which are not adequately captured by traditional benchmarks. The findings highlight a trade-off between the quality and diversity of generated samples, particularly when models are fine-tuned on instruction dataset or with human feedback. This work extends the toolkit for distribution-based NLP evaluation, offering insights into the practical capabilities and challenges that current LLMs face in generating diverse and high-quality text. We release our code and data [^1] https://github.com/AlexVerine/pr-4-llm.
@@
Florian Le Bronnec∗,1,2  Alexandre Verine∗,1 Benjamin Negrevergne1Yann Chevaleyre1Alexandre Allauzen1 1Miles, Université Paris-Dauphine, Université PSL, CNRS, LAMSADE, 75016 Paris, France 2Sorbonne Université, CNRS, ISIR, F-75005 Paris, France
@@ 1 Introduction
In recent years and months, there has been a rapid democratization of Large Language Models (LLMs), exemplified by platforms such as ChatGPT (OpenAI et al., 2024) and HuggingChat (Hugging Face, 2024). These models are now widely accessible for a diverse range of tasks, from composing emails and application letters to performing multi-document summarization, generating medical prescriptions, and even crafting poetry and novels (Hosseini et al., 2023; Skjuve et al., 2023). The expanding spectrum of applications underscores the ubiquity of LLMs in our daily lives, necessitating the development of novel evaluation frameworks to accommodate their growing relevance.
Figure 1: Precision and Recall of various models on generating the WebText dataset, with the 2 standard deviation error ellipsis. Chat and pre-trained models different behaviors are clearly captured by oevidenced.
@@
In contrast, the recent line of work on distribution-based metrics greatly departs from the sampled based evaluation. By considering LLMs and datasets as empirical distributions, the new metrics attempt to quantify how they differ and how they overlap. This shift drastically changes the scope of the evaluation. Beyond performance measures based on human references, the objective is to estimate the disparity between some data distribution ($P$) exhibited by human-written texts and the distribution learned by a LLM ($Q$), eliminating the need for aligned corpora. As exemplified by the development of MAUVE Pillutla et al. (2021), this kind of approach opens new perspectives for comparing LLMs in terms of their generative abilities rather than for some peculiar tasks.
However, many factors take part in assessing open-ended text generation and focusing on a single measure may restrict the significance of the evaluation. This is well known in the field of Information Retrieval, for example, where Precision and Recall are at the core of evaluation. Although the usage of Precision and Recall for generative models has been previously used in image generation by Sajjadi et al. (2018) and Kynkäänniemi et al. (2019), we propose to adapt them to LLMs. With an extensive set of experiments involving prominent LLMs such as Llama-2 (Touvron et al., 2023) and Mistral (Jiang et al., 2023), we show that these two new metrics significantly improve the evaluation with a better understanding and characterization of the flaws in text generation. Precision and Recall allow us to clearly distinguish between samples quality or adequacy (for Precision) and a lack of diversity in the model outputs (for Recall). Empirical results show that these two measures are necessary for a in-depth comparison of LLMs. For instance, we are able to quantify that fine-tuning models on instruction sets with human feedback significantly improves sample quality, albeit at the expense of sample diversity, as evidenced in Figure 1 by the trade-off between Precision and Recall.
@@
  • Our investigation sheds light on the impact of fine-tuning LLMs as chat models. We present empirical evidence that, while such fine-tuning can improve the quality of generated samples, it also reduce their diversity, highlighting a crucial trade-off between Precision and Recall.
@@
Figure 2: Illustration of a simple case where using MAUVE score alone fails to provide a fine-grained evaluation of quality and diversity. We consider a reference dataset $P$ composed of articles from 2 labels, World and Sport. $Q_{2}$ is made of articles from the same distribution. We compare it with two other datasets: $Q_{1}$ composed only of World articles and $Q_{3}$ composed of even numbers of World, Sport, Business and Sci/Tech articles. Relatively to $Q_{2}$, the MAUVE scores of $Q_{1}$ and $Q_{3}$ are almost identical, while Precision and Recall help differentiating how the distributions actually differ from the reference $P$.
@@ 2 Related works
Historically, assessing text generation tasks has been dependent on human resources, such as annotated texts or aligned corpora. This is particularly true in areas like machine translation and automatic summarization, where N-gram based metrics have been developed, notably BLEU (Papineni et al., 2001) and ROUGE (Lin, 2004). These metrics perform a rough comparison between the generated text and human-produced and aligned reference texts. More recent developments have introduced model-based metrics, such as BERTScore [^2] BERTScore uses precision and recall scores that differ significantly from ours, as they evaluate the similarity between a single generated text and its aligned reference. (Zhang* et al., 2020), which aim to capture semantic similarities more effectively. However, these evaluations still largely depend on "aligned references," meaning human-provided examples that constrain the expected range of generated outputs. The challenge becomes significantly greater with open-ended generation tasks, where the range of acceptable responses is vast and cannot be encapsulated by a single reference.
@@ 2.1 Standard evaluation of generative models
Perplexity is the historical measure for language modeling Bahl et al. (1977). Easy to use and cheap to compute, practical motivations clearly explain its persistence for evaluation, despite its notable limitations. As demonstrated in the work of Pillutla et al. (2021), perplexity works at the token level, only considering the surface forms, hence missing semantic information. Moreover, the metric is calculated on a per-sample basis, thereby failing to capture the diversity of generated texts.
Nowadays, the prevalent method for evaluating recent LLMs involves benchmarks with many tasks with easily verifiable correct answers (Touvron et al., 2023; Google, 2023; Beeching et al., 2023). Typically, this involves the generation of responses to short questions prompted in the instructions. The responses are then compared to a reference using metrics such as Exact-Match. While effective in assessing the comprehension and reasoning capabilities of LLMs, this kind of approach does not assess the generation skills of the models.
To overcome the limitations of automatic metrics, another trend considers human evaluation (Celikyilmaz et al., 2021), especially for complex tasks. Beyond its cost, the guidelines must be carefully designed, since various criteria could be necessary to rate texts and compare model outputs. In the end, human evaluation is prone to high variance and is difficult to reproduce. Another issue is that diversity is not evaluated is standard human evaluation protocols, since texts are rated individually or by pairs.
@@ 2.2 Evaluation of open-ended generation
Although the prevailing evaluations focus primarily on the quality of generated texts, some studies aim to characterize the diversity of these outputs. An approach involves computing the proportion of distinct N-grams generated, as demonstrated in Li et al. (2016). Another notable proposition by Zhu et al. (2018) introduces Self-BLEU which computes the BLEU score of each generated text against all other generated texts as a reference. However, N-gram matching cannot capture semantic features. For instance, a model generating random words achieves a perfect Self-BLEU score.
2.3 Quality/diversity tradeoff
The tradeoff between quality and diversity has been studied by Go et al. (2023) and Kirk et al. (2024) specifically during the human feedback alignment phase of LLMs. They draw their conclusions through traditional metrics and specifically through the summarization task. By introducing new metrics for such evaluation and focusing on extensive open-ended tasks, our methods bring new tools inspired by past research in image generation and backed by a solid theoretical framework. Our findings extend throughout the human feedback alignment phase to include both instruction tuned and pre-trained, highlighted significant differences between those types of models.
In the current landscape, most of the existing methods prioritize quality or rely on a simplistic characterization of diversity. Nevertheless, the recent introduction of MAUVE (Pillutla et al., 2021) makes a significant shift toward diversity. Drawing inspiration from the domain of image generation, this new metric compares generated texts to a reference distribution without requiring aligned texts. MAUVE consists of a divergence-based metric that captures certain properties of text and exhibits correlation with human judgment. Subsequent research by Pimentel et al. (2023) confirms the validity of this approach by leveraging common divergence measures. MAUVE and subsequent works propose a simple metric that allows us to summarize the evaluation of both quality and diversity in a single measure. Our work builds upon these efforts; we aim to demonstrate that using two distinct measures enables us to distinguish between lack of quality and lack of diversity. This distinction becomes necessary to enable a deeper and more precise understanding of generative models, as illustrated by the simple example of Figure 2.
@@ 2.3 Quality/diversity tradeoff
Figure 3: Example of distribution of images. $P$ is the reference distribution of images of the CelebA dataset Liu et al. (2015), $Q_{1}$ and $Q_{2}$ are two different distributions of images. $Q_{1}$ has high quality, but low diversity, while $Q_{2}$ has high diversity and low quality. Numbers and images are from Kynkäänniemi et al. (2019).
@@ 3 Background on Precision and Recall
Nowadays generative models for images are capable of producing ultra-realistic samples, making classical metrics such as Fréchet Inception Distance (FID) (Heusel et al., 2017) too coarse to assess the full distributional properties of the models. This jump of performance has motivated the definition of precision and diversity for generative models. This need for a finer evaluation is exemplified on Figure 3. $Q_{1}$ represents a model that produces realistic images but lacks diversity in the generated people’s appearance. Conversely, $Q_{2}$ depicts a model generating a broad range of people with lower quality. Capturing this significant tradeoff is essential for comprehending generative model behavior. Inspired by the Precision and Recall metrics of binary classifications, the pioneering work of Sajjadi et al. (2018) proposed to redefine these two terms to assess the quality and diversity of generative models. These metrics are based on a discretized estimation of the reference and model’s distribution. Kynkäänniemi et al. (2019) pointed out some limitations of those metrics and proposed a new way to compute Precision and Recall, leveraging more robust support estimation. This method is now widely adopted in the literature and is typically used along with the FID. Still, these metrics have some limitations, like the sensibility to outliers (Naeem et al., 2020). This motivates further ongoing work on this topic, including theoretical contributions (Simon et al., 2019; Naeem et al., 2020; Alaa et al., 2022; Cheema and Urner, 2023; Verine et al., 2023) or practical methods (Kim et al., 2023).
@@
Given its importance in the field of image generation, we adopt the definition of Precision and Recall introduced by Kynkäänniemi et al. (2019). Alternative definitions and approaches are discussed in Appendix A.1.
@@
Since Precision and Recall involve an estimation of the support, standard pipelines involve a projection of the samples into an latent space, using pre-trained models such as Inception-v3 (Szegedy et al., 2015) or VGG (Simonyan and Zisserman, 2015). Supports of the distributions are then estimated using a $k$-nearest neighbors algorithm. We present this algorithm in more detail in Section 4.
@@
Although using metrics characterizing quality and diversity at a distribution level is well-established in image generation (Dhariwal and Nichol, 2021; Sauer et al., 2022; Song et al., 2023), it is a very recent practice in the field of text generation. Pillutla et al. (2021) led the way in this field by introducing MAUVE, a pioneering metric that builds on the theoretical concepts of quality and diversity, as initially proposed by Djolonga et al. (2020), and detailed in Appendix A.1.
The authors showcased the intriguing characteristics of their metrics across various GPT-2 models and decoding algorithms. This metric is highly effective in evaluating the degree to which the generated distribution aligns with the reference. The methodology behind it was subsequently validated and simplified by Pimentel et al. (2023) and further evaluated with alternative frameworks by Pillutla et al. (2023).
Despite the progress made, these approaches present a challenge: interpreting them as a measure of quality and diversity is not straightforward, as they condense these two concepts into a single divergence measure. Characterizing Precision and Recall independently has only been mentioned by Pillutla et al. (2023) who conducted some toys experiments on small models or decoding algorithms.
Therefore, we believe that Precision and Recall for text generation are in its early stages. We argue that a deeper understanding of these metrics and their practical implications is needed.
@@ 4 Precision and Recall to assess text generation
In this paper, we introduce a framework for calculating Precision and Recall between distributions on texts. These distributions characterize a dataset, or a model, and its generative capability. The extended version of Precision and Recall builds upon the notion of support of distributions, which is in our case challenging to define and estimate.
@@ 4.1 Pipeline to compute Precision and Recall
    Sampling. Select $N$ sequences randomly in the dataset to build the set of references: $\mathcal{X}^{\mathrm{ref}}=\left\{{\bm{x}}_{1}^{\mathrm{ref}},\dots,{\bm{x}}_{N}^{\mathrm{ref}}\right\}.$ Sample $N$ sequences from the distribution $Q$ to assess and build the set of outputs: $\mathcal{X}^{\mathrm{out}}=\left\{{\bm{x}}_{1}^{\mathrm{out}},\dots,{\bm{x}}_{N}^{\mathrm{out}}\right\}.$
@@
    Precision and Recall computation.
@@
To choose the other parameters – the number of samples $N$ and the neighbor parameter $k$ – we investigate how Precision and Recall are affected by their variations. We use the WebText generation task with Llama-2 7B Chat (comprehensively described in Section 5.3) as a test case, prompting the model to generate between 100 and 10,000 samples. The findings are depicted in Figure 6. Additional details can be found in Section B.4.
Our observations are as follows: (1) Precision and Recall values stabilize with an increasing $N$, plateauing at approximately $N=3000$. (2) As $k$ increases, Precision and Recall values gradually approach 1, aligning with the expectation that a larger $k$ leads to a wider estimated support. In the rest of the paper, we set $N$ to $4000$ samples and $k$ to $4$.